# LayoutCoT: Unleashing the Deep Reasoning Potential of Large Language Models for Layout Generation

### v1

@@ 4.5 Results on Constraint-Explicit Layout Generation
Table 3 presents the results of the quantitative evaluation, where we compare LayoutCoT with several training-based baselines and the training-free method LayoutPrompter. From the metrics, it is evident that, even without any training, LayoutCoT achieves optimal performance on most indicators. The corresponding qualitative results are shown in Figure. 5, which demonstrates that LayoutCoT exhibits outstanding layout control capabilities and generation quality. LayoutCoT not only effectively meets various input constraints—including element type constraints, size constraints, and element relationship constraints—but also produces visually pleasing layouts by avoiding overlap between elements and enhancing overall alignment. These findings collectively validate the effectiveness of LayoutCoT.
@@ 4.6 Results on Text-to-Layout
The quantitative metrics are presented in Table 4. Because text-to-layout is one of the most challenging tasks in layout generation, LayoutCoT falls short in the mIoU metric compared to the parse-then-place [27] method. However, it performs exceptionally well on the remaining three metrics. In many scenarios, LayoutCoT more effectively fulfills textual descriptions and excels in aesthetics, avoiding element overlap and maintaining element alignment.
Table 4: Quantitative comparison with baselines on text-to-layout. LayoutCoT denotes the results when using ’DeepSeek-R1’ as the model with the CoT module removed.
Methods mIoU $\uparrow$ FID $\downarrow$ Align $\downarrow$ Overlap $\downarrow$
Mockup [15] 0.193 37.012 0.0059 0.4348
parse-then-place [27] 0.684 2.959 0.0008 0.1380
@@ 4.7 Ablation Studies
To demonstrate the effectiveness of our approach, we conducted an overall ablation study, as well as ablations focusing on the selection method for prompting examples in the Layout-aware RAG and on the choice of the number of RAG layouts $k$ in the CoT module. All our ablation studies are conducted on the PKU dataset within the representative Content-Aware layout generation task.
Overall ablation study. In Table. 5, we demonstrate the role of each module. It is evident that the Layout-aware RAG and the CoT module complement each other, yielding excellent results across all metrics. Regarding the higher Overlap score when the CoT module is used, we provide an explanation in Sec. 4.4. In reality, although the Overlap metric increases, the CoT module effectively rectifies the layout’s unreasonableness and lack of aesthetic appeal.
Table 5: Abalation study of each method.
@@
0.251 0.171 0.237 0.0021 0.824 0.809 0.0005 0.997
0.209 0.161 0.231 0.0004 0.931 0.844 0.0014 0.999
0.207 0.074 0.289 0.0002 0.980 0.946 0.0013 1.000
The Method of the Layout-aware RAG. Table. 6 reports the experimental results obtained with different RAG‑based strategies for selecting prompting examples. Specifically, random denotes randomly selecting $k$ examples; the mIoU strategy chooses examples according to the maximum IoU between saliency maps; DreamSim [8] ranks candidates by their DreamSim scores. The results show that random selection degrades layout‑generation performance—sometimes performing even worse than omitting RAG entirely. By contrast, the mIoU strategy, which assesses relevance between saliency maps, offers a modest improvement, while DreamSim—considering both image and layout coherence—yields further gains over mIoU. Finally, the ltsim method, which measures overall layout similarity, proves the most effective and achieves the best generation results.
Table 6: Abalation study of Layout-aware RAG.
[…]
- 0.209 0.161 0.231 0.0004 0.931 0.844 0.0014 0.999
@@
DreamSim [8] 0.252 0.086 0.273 0.0004 0.972 0.932 0.0008 0.999
ltsim 0.207 0.074 0.289 0.0002 0.980 0.946 0.0013 1.000
Selection of $k$. In Table. 7, we present the impact of different numbers of prompting examples on the CoT module’s performance. It can be observed that when $k$=2, the results are relatively poor. This is because having fewer prompting examples means the LLM lacks sufficient data references for deep reasoning, thus impairing its inference capability. On the other hand, when $k$=6, the outcome is essentially the same as when $k$=4. We think that once the number of prompting examples reaches a certain threshold, increasing $k$ further does not significantly stimulate deeper reasoning in the LLM, and also consumes more resources. Therefore, we choose $k$=4.
Table 7: Abalation study of the selection of $k$.
[…]
2 0.232 0.119 0.282 0.0009 0.835 0.797 0.0019 0.999
4 0.207 0.074 0.289 0.0002 0.980 0.946 0.0013 1.000
6 0.216 0.075 0.301 0.0004 0.968 0.939 0.0010 1.000
Effectiveness of the Multi‑Stage of CoT. To verify that a multi‑stage chain of thought is more effective than a single‑stage approach, we combined all sub‑problems from the multi‑stage process into a single question and performed a one‑stage refinement. The quantitative metrics, presented in Table. 8, show that the single‑stage strategy performs considerably worse than the multi‑stage variant. This disparity arises because layout generation follows a logical sequence: resolving sub‑problems step by step enhances layout rationality, whereas tackling multiple issues simultaneously is impractical and yields limited improvements.
Table 8: Abalation study of the effectiveness of multi-stage.
[…]
1 0.219 0.167 0.241 0.0005 0.933 0.851 0.0017 0.999
3 0.207 0.074 0.289 0.0002 0.980 0.946 0.0013 1.000

### v3

@@ 4.5 Results on Constraint-Explicit Layout Generation
Table 3 presents the quantitative evaluation comparing LayoutCoT with several training-based baselines and the training-free method LayoutPrompter. From the metrics, it is evident that, even without any training, LayoutCoT achieves optimal performance on most indicators. The corresponding qualitative results in Figure 5 demonstrate that LayoutCoT exhibits outstanding layout control capabilities and generation quality. LayoutCoT not only effectively meets various input constraints–including element type, size, and relationship constraints–but also produces visually pleasing layouts by avoiding overlap between elements and enhancing overall alignment. These findings validate the effectiveness of LayoutCoT.
@@ 4.6 Results on Text-to-Layout
The quantitative metrics are presented in Table 4. Because text-to-layout is one of the most challenging tasks in layout generation, LayoutCoT falls short in the mIoU metric compared to the parse-then-place Lin et al. (2023a) method. However, it performs exceptionally well on the remaining three metrics. In many scenarios, LayoutCoT more effectively fulfills textual descriptions and excels in aesthetics, avoiding element overlap and maintaining element alignment.
Table 4: Quantitative comparison with baselines on text-to-layout. We compare our method with the baselines Mockup Huang et al. (2021), parse-then-place Lin et al. (2023a), and LayoutPrompter.
Methods mIoU $\uparrow$ FID $\downarrow$ Align. $\downarrow$ Overlap $\downarrow$
Mockup 0.193 37.012 0.0059 0.4348
parse-then-place 0.684 2.959 0.0008 0.1380
@@ 4.7 Ablation Studies
We validate the effectiveness of each component on the PKU dataset, as summarized in Table 5.
Impact of Modules. Table 5 (a) demonstrates that both the Layout-aware RAG and CoT modules individually improve performance over the baseline. Specifically, RAG enhances structural alignment (Align $0.0021\!\rightarrow\!0.0005$), while CoT effectively reduces occlusion. The full LayoutCoT framework achieves the best balance across all metrics, confirming the complementarity of retrieval context and reasoning capabilities. The slight rise in Overlap is a necessary trade-off to correct the baseline’s tendency to generate unrealistically small, crowded boxes.
Retrieval Strategy Analysis. Regarding the RAG module, Table 5 (b) shows that our LTSim-based retrieval significantly outperforms Random, mIoU, and DreamSim strategies by effectively capturing holistic layout similarity. Furthermore, we investigate the number of retrieved examples $k$ in Table 5 (c), finding that $k=4$ offers the optimal trade-off between providing sufficient context and minimizing noise.
CoT Design Validation. Finally, we analyze the reasoning mechanism. Table 5 (d) reveals that our multi-stage strategy markedly surpasses the single-stage approach, validating the necessity of decomposing complex layout tasks into logical steps. Moreover, Table 5 (e) confirms that our domain-specific manually designed CoT outperforms the generic Self-Refine mechanism in 7 out of 8 metrics, ensuring superior layout rationality.
Table 5: Comprehensive Ablation Studies on the PKU dataset. We systematically analyze the impact of (a) overall modules, (b) layout-aware retrieval strategies, (c) the number of retrieved examples ($k$), (d) reasoning stages, and (e) refinement methodologies. The best results are highlighted in bold.
@@
(a) Overall Module Ablation
Baseline (No RAG/CoT) 0.251 0.171 0.237 0.0021 0.824 0.809 0.0005 0.997
w/ RAG only 0.219 0.167 0.241 0.0005 0.933 0.851 0.0017 0.999
w/ CoT only 0.209 0.161 0.231 0.0004 0.931 0.844 0.0014 0.999
LayoutCoT (Full) 0.207 0.074 0.289 0.0002 0.980 0.946 0.0013 1.000
(b) Layout-aware RAG Strategy
No RAG 0.209 0.161 0.231 0.0004 0.931 0.844 0.0014 0.999
@@
DreamSim 0.252 0.086 0.273 0.0004 0.972 0.932 0.0008 0.999
LTSim (Ours) 0.207 0.074 0.289 0.0002 0.980 0.946 0.0013 1.000
(c) Number of Retrieved Examples ($k$)
$k=2$ 0.232 0.119 0.282 0.0009 0.835 0.797 0.0019 0.999
$k=4$ (Ours) 0.207 0.074 0.289 0.0002 0.980 0.946 0.0013 1.000
$k=6$ 0.216 0.075 0.301 0.0004 0.968 0.939 0.0010 1.000
(d) CoT Reasoning Stages
Single-Stage (1) 0.219 0.167 0.241 0.0005 0.933 0.851 0.0017 0.999
Multi-Stage (3) 0.207 0.074 0.289 0.0002 0.980 0.946 0.0013 1.000
(e) Refinement Method
Self-Refine 0.153 0.149 0.213 0.0012 0.868 0.806 0.0029 1.000
Manual CoT (Ours) 0.207 0.074 0.289 0.0002 0.980 0.946 0.0013 1.000
@@ 5 Conclusion
6 Limitations
Our work focuses on unlocking the reasoning potential of LLMs, but this comes with trade-offs. The iterative nature of our CoT mechanism, while boosting performance, incurs higher computational overhead than non-autoregressive models. Additionally, since our method leverages in-context learning, its upper bound is influenced by the retrieval database; expanding the database with more diverse layouts would be necessary to handle extremely rare or complex constraints effectively.
Appendix A Prompt Examples of LayoutCoT
A.1 Content Aware Layout Generation
Stage 1: Initial Adjustment
System Prompt Role. You are an advanced layout design assistant, LayoutGPT. We will refine the layout in 3 separate steps. Each step is isolated, and you only have the references provided in the current context. Goal. Adjust element positions to achieve reasonable alignment without overlap. Procedure. 1. Adjust only text and logo elements (modify x/y position, size, and font-size). 2. Keep object and underlay locked (do not move or alter them). 3. Avoid overlap between text, logo, and object — even 1px overlap is disallowed. 4. Achieve reasonable alignment: for each logical group, logo should ideally appear above the corresponding text. 5. Return only updated HTML content with no extra explanation. Note. You may learn spatial relationships and layout logic from the reference layouts, but you must not directly copy element coordinates.
User Prompt Below are $LEN_{TOPK}$ reference layouts: {{TOPK_HTML_STR}} Here is the initial layout needing refinement: {{CURRENT_HTML}}
Stage 2: Underlay Refinement
System Prompt Role. You are LayoutGPT. Continue refining the layout. Goal. Optimize the background elements based on the locked foreground. Procedure. 1. Keep object, text, and logo elements exactly as they were after Stage 1 (locked). 2. Adjust only underlay elements (background rectangles). 3. Avoid unintended overlap: underlay must not intersect with unrelated content. 4. Improve layout metrics: metrics_und_l: If underlay > target, maximize IoU with associated text/logo. metrics_und_s: If underlay $\leq$ target, ensure it fully contains the corresponding text/logo. 5. Return only updated HTML content.
User Prompt Below are the $LEN_{TOPK}$ reference layouts: {{TOPK_HTML_STR}} Here is the layout after Stage 1: {{STAGE_1_HTML}} Please do a final check and minimal fixes as needed. Return only the final HTML.
Stage 3: Final Check
System Prompt Role. You are LayoutGPT. Perform the final check. Goal. Ensure zero unintended overlap and perfect alignment. Procedure. 1. Check for unintended overlap across all elements (text, logo, underlay, object). Ensure underlay only overlaps its own group. 2. Ensure alignment is preserved (e.g., logo above text). 3. Perform minimal fixes for small issues (e.g., slight 1-2px overlap, container issues). 4. Return only the final HTML result.
User Prompt Below are the $LEN_{TOPK}$ reference layouts: {{TOPK_HTML_STR}} Here is the layout after Stage 2: {{STAGE_2_HTML}} Please do a final check and minimal fixes as needed. Return only the final HTML.
A.2 Constraint-Explicit Layout Generation (Gen-T)
We provide the prompt for the Gen-T task on the PubLayNet dataset as a representative example.
Stage 1: Main Element Placement
System Prompt Context. We have 5 element types: 1) text, 2) title, 3) list, 4) table, 5) figure. Goal. Initial placement of core text elements. Procedure. 1. Move/resize only title (2) and text (1). 2. Keep list (3), table (4), and figure (5) locked. 3. Minimize overlap between title & text and place them with good alignment. 4. Keep FID & maxIoU in mind. 5. Return only updated HTML.
User Prompt Here are $LEN_{TOPK}$ reference layouts: {{TOPK_HTML_STR}} Initial layout needing Stage 1 refinement: {{CURRENT_HTML}} Please adjust only title(2) & text(1) to reduce overlap and improve alignment.
Stage 2: Complex Element Placement
System Prompt Goal. Place complex elements while respecting Stage 1. Procedure. 1. Place/resize list (3), table (4), and figure (5). 2. Keep title (2) & text (1) from Stage 1 locked if possible. 3. Further reduce overlap among all 5 elements. 4. Refine alignment (e.g., align edges of tables with text). 5. Begin improving FID and maxIoU. 6. Return only updated HTML.
User Prompt $LEN_{TOPK}$ reference layouts: {{TOPK_HTML_STR}} Layout after Stage 1: {{STAGE_1_HTML}} Please refine list(3), table(4), figure(5) now. Return updated HTML.
Stage 3: Global Optimization
System Prompt Goal. Final thorough check for metrics optimization. Objectives. Alignment: Lower score implies better alignment. Overlap: Lower score implies minimal collisions. FID: Reduce unnatural layout styles. maxIoU: Match reference bounding arrangements. Procedure. 1. Enumerate each pair for overlap. 2. Check alignment lines or columns. 3. Evaluate if FID can be lowered by slight repositioning. 4. Shift/resize elements to improve IoU with references. 5. Provide brief chain-of-thought reasoning, then output final HTML.
User Prompt $LEN_{TOPK}$ reference layouts: {{TOPK_HTML_STR}} Layout after Stage 2: {{STAGE_2_HTML}} Perform your step-by-step reasoning to finalize alignment, overlap, FID, maxIoU. Return final HTML.
A.3 Text-to-Layout Generation
Stage 1: Parsing and Rough Layout
System Prompt Role. You are LayoutGPT. Goal. Parse text and establish initial structure. Procedure. 1. Parse text to identify required elements; add missing or remove extraneous ones. 2. Attempt strong overlap reduction from the start (moves or resizes). 3. Perform basic alignment ($\pm$10px). 4. Partially match reference bounding arrangement for mIoU if feasible. 5. Return updated HTML only.
User Prompt References (topk) total $LEN_{TOPK}$: {{REFERENCES_STR}} Current layout: {{CURRENT_HTML}} Text describing needed elements: {{TEXT_DESCRIPTION}} Please add/remove elements per text, reduce overlap, and do basic alignment. Return updated HTML.
Stage 2: Overlap Resolution
System Prompt Goal. Aggressive overlap reduction. Procedure. 1. Reduce Overlap to near zero (Top Priority). 2. Refine alignment lines ($\pm$5px). 3. Match bounding arrangement from references for higher mIoU. 4. Add or remove minor elements if beneficial. 5. Return updated HTML only.
User Prompt References: {{REFERENCES_STR}} Layout after Stage 1: {{STAGE_1_HTML}} Focus on Overlap $\to$ 0, align row/col $\pm$5px. Return updated HTML.
Stage 3: Semantic-Agnostic Refinement
System Prompt Goal. Final geometric polish. Procedure. 1. Temporarily ignore text meaning. Force Overlap = 0 among main elements. 2. Fix leftover collisions by moves/resizes. 3. Push bounding arrangement for better mIoU. 4. Fix FID style if it looks awkward. 5. Return updated HTML only.
User Prompt References: {{REFERENCES_STR}} Layout after Stage 2: {{STAGE_2_HTML}} Ignore text semantics, finalize Overlap=0, fix FID. Return updated HTML.